# Path-based packing of icosahedral shells into multi-component aggregates

### v2

@@
Path-based packing of icosahedral shells into multi-component aggregates
Nicolò Canestrari, Diana Nelli, Riccardo Ferrando
Dipartimento di Fisica, Università di Genova, Via Dodecaneso 33, 16146 Genova, Italy
To whom correspondence should be addressed;
E-mail: diana.nelli@edu.unige.it; riccardo.ferrando@unige.it
@@ Introduction
The research on multi-components nanoaggregates is extremely active and spans many different fields. High-entropy alloy nanocrystals, consisting of nanometer-sized solid solutions of five or more elements, have attracted much attention due to their enhanced structural stability and catalytic activity [1, 2, 3, 4]. Ordered architectures are explored as well. Among them, the assembly of concentric shells of different compositions into multilayer aggregates is a widely employed tool to protect or functionalize the core [5, 6], improve the stability [7, 8, 9, 10], and adjust the surface properties [11, 12] of the nanostructure.
The wide range of possible compositions and the extremely rich configurational space of multi-component systems are key to their broad success. However, such inherent complexity poses major challenges to the design and synthesis of nanoaggregates with well-defined and durable configurations. A general theory for the prediction of stable multicomponent structures would be an essential reference for the design and synthesis of nanoparticles for customised applications. However, such a theory is lacking at present.
@@
The key point of our theoretical approach is the mapping of multi-shell structures into paths in the hexagonal lattice. The mapping naturally leads to the design of new multi-shell sequences corresponding to new series of icosahedral magic numbers. Furthermore, the mapping allows us to predict which sequences exhibit spontaneous symmetry breaking from achiral to chiral structures. In the field of metal clusters and nanoparticles, chiral icosahedra are especially interesting for their applications to catalysis [41]. In all cases, we demonstrate that icosahedra are stabilized by the size mismatch between particles of different shells and evaluate the optimal mismatch for energetic stability. We note that, due to the contraction of pair distances between particles in adjacent shells [13], the icosahedron is naturally suitable for accommodating particles of different sizes in different shells. Our approach is generalized also to anti-Mackay shells, so that it allows us to derive a design principle for the structures observed in colloidal aggregates [25, 26, 27], and to predict new ones.
Our design strategy is validated by numerical calculations for several model systems and by ab initio calculations for alkali metal clusters. The predictions of our theory are further confirmed by simulations of the growth of alkali and transition metal nanoparticles including up to four different elements. These simulations show that atoms naturally self-assemble into the predicted multi-shell icosahedral structures, including those with symmetry-breaking shell sequences. Although the applications presented below concern atomic clusters and nanoparticles, due to the general character of its basic assumptions, our theory can be applied to the design of aggregates of other particle types, such aggregates of colloidal particles and complex molecules of biological interest.
@@
Here we develop our theoretical framework, whose starting point of is the well-known approach of Caspar and Klug [14, 16, 15], which was originally proposed for rationalizing and predicting the architecture of icosahedral viral capsids. Specifically, Caspar and Klug (CK) developed a general method for the construction of individual icosahedral shells, which is based on cutting and folding leaflets from the two-dimensional hexagonal lattice and produces achiral and chiral arrangements of particles on the icosahedral surface.
@@
In the CK theory, the triangulation number $T$ of an icosahedral shell is defined as the square length of the triangular edge, and is calculated as $T=h^{2}+k^{2}+hk$. Edge and radius of the shell are $\sqrt{T}$ and $\sin(2\pi/5)\sqrt{T}$, respectively. Assigning one particle to each lattice point, the shell contains $210\,T+$ 2 particles (see Supplementary Note 1.1).
Shells are achiral or chiral depending on the angle $\theta_{CK}$ between the $h$-axis and the segment of the CK construction. Achiral shells correspond to segments with $\theta_{CK}=0^{\circ},30^{\circ}$ and 60. Segments on a coordinate axis correspond to the achiral shells described by Mackay (MC) in his work on the packing of equal spheres [13]. Shells built on segments on the diagonal ($h=k$, $\theta_{CK}=30^{\circ}$) are here called of Bergman (BG) type, since the smallest is the outer shell of the Bergman cluster [30, 38]. All other shells are chiral, with enantiomers symmetrically placed with respect to the diagonal. We remark that, in the CK theory, an icosahedral shell is uniquely determined by the segment endpoint $(h,k)$ in the hexagonal lattice; therefore, in the following, icosahedral shells will be denoted by their $(h,k)$.
[…]
Figure 2: Mapping icosahedra into paths. a Path of a Mackay icosahedron: $i=5$ shells are assembled on top of each other. b The three possible cases when choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$ at each step of the path.
@@
In general, paths can be drawn by allowing different increments at each step. Here we deal with the simplest generalization, which consists of choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$. In this case, the number of shells in a path from the origin to a point $(h,k)$ is $i=h+k+1$. Such paths are inspired by the Mackay path of Figure 2a, in which one index is incremented at steps of one, whereas the other is kept constant and equal to zero. Allowing for different elementary steps in the hexagonal plane allows to build a wide variety of icosahedral structures, which retain the densely-packed character of the Mackay icosahedron.
@@
When assembling shells in physical systems, one must bear in mind that in the icosahedron the radius is shorter than the edge by $\sin(2\pi/5)\simeq 0.9511$, which has a direct effect on the packing of concentric shells. In the Mackay icosahedron of equal spheres [13], the distance between spheres in neighbouring shells is shorter by about $5\%$ than that between spheres in the same shell. Similar considerations hold for shells belonging to other chirality classes, assembled according to the path rules identified so far; in some cases, the difference between intra-shell and inter-shell nearest-neighbour distances is even larger. These considerations naturally lead to the idea of assembling shells in which particles in different shells have different sizes, which we better clarify below.
We consider the case of a core with $i$ shells, to which we add the outer shell $i+1$. Particles in the core and in the outer shell have different sizes. We define the size mismatch $sm_{i,i+1}=(d_{i+1}-d_{i})/d_{i}$, with $d_{i+1}$ and $d_{i}$ particle sizes in shells $i+1$ and $i$, respectively.
Figure 3: Optimal size mismatch. a Binding energy per atom for clusters made of a core with five MC shells plus a sixth shell, which is either MC or Ch1. Interactions are of Lennard-Jones (LJ) type (see Methods Section). The particles of the sixth shell differ from those in the core only by their size. The energy is in units of the $\varepsilon$ of the LJ potential. The structures are shown in b. c Comparison of the optimal mismatch between steps $(0,k)\to(0,k+1)$ and $(0,k)\to(1,k)$, corresponding to additional MC and Ch1 shells on an MC core, respectively. The optimal mismatch of Eq. 3 (GEO values) is compared to Lennard-Jones (LJ) and Morse potential results. The Morse potential data are given for three values (4,5,6) of the parameter $\alpha$, that regulates the width of the potential well (see Methods Section). d Comparison between steps $(1,k)\to(1,k+1)$ and $(1,k)\to(2,k)$, corresponding to additional Ch1 and Ch2 shells on a core containing $k$ MC and one Ch1 shells, respectively. The optimal mismatch of Eq. 3 (GEO) is compared to LJ data
In the example of Figure 3a-b, the core is made of five MC shells, i.e. it is terminated by the $(0,4)$ shell. According to our path rules, the outer shell can be either the $(0,5)$ MC shell (same class) or the $(1,4)$ Ch1 shell (different class). We consider a simple model, in which all particles interact by the well-known Lennard-Jones (LJ) potential; specifically, the interaction energy is the same for all particles, the only difference being the equilibrium distance of pair interactions, which accounts for particles of different sizes (see the Methods section in the supplementary materials). In Figure 2a, we calculate the binding energy depending on the size mismatch, and determine the mismatch that minimizes the energy of the whole aggregate. For both MC and Ch1 shells this optimal mismatch is positive, i.e. it is favourable to have bigger particles in the outer shell than in the core, and it is larger for Ch1 than for MC shells.
The optimal mismatch can be estimated also by geometric packing arguments, which are discussed in details in Supplementary Note 1.4. Here we recall only the main assumption and give the final result. In order to evaluate the optimal mismatch between particles of shells $i$ and $i+1$ it is reasonable to impose that
@@
where $r_{i}$ and $r_{i+1}$ are the radii of the respective shells, $d_{i}$ and $d_{i+1}$ are the particle diameters in these shells. For atomic systems, these diameters are estimated from the nearest-neighbour distance of atoms in their crystal lattice. Recalling that the relation between the radii and the triangulation numbers of the shells is
$r_{i}=d_{i}\sin\left(\frac{2\pi}{5}\right)\sqrt{T_{i}}$ (2)
@@
$sm_{i,i+1}=\frac{2\sin\left(\frac{2\pi}{5}\right)(1+\xi)\sqrt{T_{i}}+1}{2\sin\left(\frac{2\pi}{5}\right)\sqrt{T_{i+1}}-1}-1.$ (3)
@@
In Figure 3c-d the results of Eq. 1 are compared to those for LJ and Morse clusters. Data in Figure 3c are obtained as in Figure 3a, but for different sizes of the Mackay core, whereas in Figure 3d we consider a cluster made of a Mackay core plus a Ch1 shell, to which we add a further shell of either Ch1 or Ch2 type. In all cases the agreement is good. Eq. 1 is thus a reliable guide for the semi-quantitative evaluation of the optimal mismatch, which demonstrates the key role of geometric factors in determining it.
@@
Our theory is now applied to the path-based design of icosahedral aggregates. In order to construct a multi-shell icosahedral aggregate, a path is drawn according to the rules in Figure 2 and, for each step, the optimal mismatch is estimated by Eq. 1. Then, particles of the appropriate sizes are associated with each shell.
Figure 4: A path with spontaneous symmetry breaking. a A path connecting all BG shells through their nearby chiral shells. Black and grey arrows indicate paths with right and left-handed chiral shells. A generic path of this type may alternate both chiralities, corresponding to combinations of black and grey arrows. Shells are enumerated along the black-arrow path. b Optimal mismatch $sm_{i,i+1}$ according to Eq. 1 for the path in a. Points connected by brown and green lines correspond to changes of species every shell and every two shells. The optimal $sm_{i,i+1}$ are compared to the mismatch between pairs of atomic species, indicated by squares and diamonds. The triangles correspond to the optimal mismatch between two MC shells and two Ch1 shells. c Structures for $i=2,3,4$ along the path of a, with their compositions related to specific systems.
An example is shown in Figure 4, where the path connecting BG shells through neighbouring chiral shells is considered (Figure 4a). This path alternates class-conserving and class-changing steps, spontaneously breaking mirror symmetries after the first BG shell. It produces a new series of magic numbers
@@
with $b=4$ and $b=9$ for even and odd $i$, i.e. $N_{i}=1,13,45,117,239,431..$ (see Supplementary Note 1.2).
The path of Fig. 4a is used to design multi-species alkali clusters. The values of size mismatch between alkali atoms, estimated from nearest-neighbour distances in bulk crystals, are in the range of the optimal values for this path (Fig. 4b). In addition, these species present a weak tendency to mix and the bigger atoms have a smaller cohesive energy; this produces a general tendency for the bigger atoms to stay in the surface layers, which, as we have seen in the previous Section, is exactly what is needed for stabilizing icosahedral aggregates.
The clusters in Figure 4c correspond to $i=3$ (Na13@K32, Na@${}_{13}@$Rb32) and to $i=4$ (Na@${}_{13}@$K@${}_{32}@$K72, Na@${}_{13}@$Rb@${}_{32}@$Rb72, Na@${}_{13}@$K@${}_{32}@$Rb72). In these clusters, the atomic species is always changed in class-changing steps, where a large mismatch is required, while it is changed or not in class-conserving steps, where the optimal mismatch is moderate, and therefore zero mismatch is expected to be acceptable. The stability of these clusters is verified by Density Functional Theory (DFT) calculations (data are reported in Table 1 and Supplementary Tables 1-3), specifically by checking that exchanges of pairs of different atoms in adjacent shells produce energy increases. In contrast, if MC shells of these species are assembled in the same order, the resulting clusters are energetically unstable with respect to exchanges of atomic pairs, because the mismatch between species is too large for MC shells. Our theory therefore provides new possibilities for constructing icosahedra in systems where the traditional Mackay icosahedra would not be stable.
3 SHELLS 4 SHELLS
MC@BG MC@MC MC@BG@Ch1 MC@MC@MC
Na@K +0.223 $-0.345$ +0.164 $-0.064$
Na@Rb +0.237 $-0.711$ +0.205 $-0.081$
Na@K@Rb -- -- +0.180 $-0.006$
Table 1: DFT data for atom pair exchanges. In the initial configurations, these clusters always contain a core with MC shells $(0,0)$ and $(0,1)$ made of Na atoms, while the other shells are made of K, Rb or K and Rb atoms. For MC@BG, the $(1,1)$ BG shell is added to the core, whereas for MC@BG@Ch1 a further $(1,2)$ shell is added (see Supplementary Figure 6a). For MC@MC, the $(0,2)$ shell is added to the core, whereas for MC@MC@MC a further $(0,3)$ shell is added. Exchanges of atom pairs are made on the initial configuration: a Na atom in the $(0,1)$ MC shell is exchanged with a neighbouring K/Rb atom of the outer shell (see Supplementary Figure 6b-e). The energy differences (in eV) between the configurations after and before the exchange of the atomic pair are reported, for the most favourable between the two possible inequivalent exchanges. The exchange is favourable for negative differences, unfavourable otherwise. Complete data are reported in Supplementary Tables 1-3.
@@ 3 SHELLS 4 SHELLS
We have demonstrated how to geometrically construct multi-shell icosahedra and verified their energetic stability in various systems. Another important point is to understand how they grow dynamically in physical processes. To this end, we have performed molecular dynamics (MD) growth simulations [44], in which atoms are deposited on pre-formed clusters. This type of simulations has been used to interpret previous nanoparticle growth experiments [24, 44]. The results are shown in Figure 5 and in Supplementary Figures 7-11.
Figure 5: Growth sequences of chiral icosahedra. a Rb atoms are deposited on an icosahedral Na@${}_{13}@$Rb32 seed. b The seed is Na@${}_{13}@$Rb104, on which Cs atoms are deposited. c Size mismatch for some transition metal pairs compared to the optimal mismatch for a MC and a Ch1 shell on a MC core (green and brown lines, respectively), as a function of core size. d-e Snapshots from growth simulations and corresponding paths for the deposition of Ag atoms on d Ni147 and g Cu561 cores. All simulation snapshots in a-b, d-e are taken at magic sizes for the corresponding paths. In the top and bottom rows of the snapshot sequences we show the cluster surface and its cross section, respectively.
As a first case, we consider the growth of multi-element alkali clusters. In the previous Section we have demonstrated the stability of non-trivial multi-shell structures, which comprise BG and chiral shells in the same aggregate. Here we check whether these structure can be grown in a physically realistic process.
@@
At this point, according to our rules, two non-equivalent steps are possible: the class-conserving step to $(1,3)$ and the class-changing step to $(2,2)$. These steps correspond to different optimal mismatches, i.e. 0.06 and 0.11, respectively, as estimated by Eq. 1. Since we are keeping on depositing Rb atoms (i.e. with zero mismatch), we expect growth to proceed by the $(1,3)$ step, which has the lower optimal mismatch. Also in this case, our prediction is verified by the growth simulation of Figure 5a, which further continues within the Ch1 class.
@@
Further confirmation of path rule predictivity is provided in Figure 5c-e. In Figure 5c, the size mismatch for the pairs AuNi, AuCo, AuFe, AgNi, AgCo, AgCu is compared with the optimal estimates obtained from Eq. 1 for the addition of one MC and one Ch1 shell on a Mackay core. The mismatch much better corresponds to Ch1 than to MC shells. For AuCo, AuFe, AuNi, AgNi and AgCo, Ch1 shells should grow on Mackay cores of 147 or 309 atoms ($k=3$ or $4$), while for AgCu the core should be larger, of 561 atoms ($k=5$). This is confirmed by MD simulations (Figure 5d-e) in which a Ch1 shell spontaneously forms when depositing Ag atoms on Ni147 or Cu561 Mackay cores. The growth continues within the Ch1 class if further Ag atoms are deposited. More results are presented and discussed in Supplementary Note 3.
@@
Finally, we note that almost perfect shell-by-shell growth is achieved in all simulations due to the fast diffusion of deposited atoms on top of the close-packed shells [24].
@@
[…]
Figure 7: Growth of generalized anti-Mackay structures. a Optimal mismatch for one AM1 shell on a Mackay core ($(0,k)\to(1,k)^{*}$) as a function of core size, compared with Ch1 ($(0,k)\to(1,k)$) and Mackay shells ($(0,k)\to(0,k+1)$), for Lennard-Jones clusters. GEO values are calculated by the formula derived in Supplementary Note 1.5. The mismatch between Ag and Cu is indicated. b,c Snapshots from MD growth simulations (with cluster surfaces and cross sections shown in top and bottom lines, respectively) and corresponding paths, with AM shells indicated by squares. In b Ag atoms are deposited on a Mackay Cu147 core. In c Pd atoms are deposited on a Ni core and then Ag atoms are deposited on the PdNi cluster. In the top and bottom rows of the snapshot sequences we show the cluster surface and its cross section, respectively.
@@
From there on, the growth proceeds in a different way from that of Figure 5e: a second shell of AM2 type is formed, followed by a shell of BG type, i.e. the class is changed at each step, even though atoms of the same type are deposited. This behaviour is due to the kinetic of the growth process. Specifically, it is due to fourfold adsorption sites on the surface of AM shells, which act as adatom traps and naturally lead to the formation of AM shells of higher class (see Supplementary Figure 11 for details on the growth mechanism). The fourfold traps make the growing structure much less smooth than in the case of Figure 5e. We note that, even though kinetic effects dominate the process, the growth proceeds step by step by incrementing only one shell index, as predicted by our rules.
@@
Outlook

### v3

# General theory for packing icosahedral shells into multi-component aggregates

@@
Abstract
@@ Introduction
The research on multi-components nanoaggregates is very active and spans many different fields. High-entropy alloy nanocrystals, consisting of nanometer-sized solid solutions of five or more elements, have attracted much attention due to their enhanced structural stability and catalytic activity [1, 2, 3, 4]. Ordered architectures are explored as well. Among them, the assembly of concentric shells of different compositions into multilayer aggregates is a widely employed tool to protect or functionalize the core [5, 6], improve the stability [7, 8, 9, 10], and adjust the surface properties [11, 12] of the nanostructure.
The wide range of possible compositions and the huge configurational space of multi-component systems are key to their broad success. However, such inherent complexity poses major challenges to the design and synthesis of nanoaggregates with well-defined and durable configurations. A general theory for the prediction of stable multicomponent structures would be an essential reference for the design and synthesis of nanoparticles for customised applications. However, such a theory is lacking at present.
@@
The key point of our theoretical approach is the mapping of multi-shell structures into paths in the hexagonal lattice. The mapping naturally leads to the design of multi-shell sequences corresponding to alternative series of icosahedral magic numbers. Furthermore, the mapping allows us to predict which sequences exhibit spontaneous symmetry breaking from achiral to chiral structures. In the field of metal clusters and nanoparticles, chiral icosahedra are especially interesting for their applications to catalysis [41]. In all cases, we demonstrate that icosahedra are stabilized by the size mismatch between particles of different shells and evaluate the optimal mismatch for energetic stability. We note that, due to the contraction of pair distances between particles in adjacent shells [13], the icosahedron is naturally suitable for accommodating particles of different sizes in different shells. Our approach is generalized also to anti-Mackay shells, so that it allows us to derive a design principle for the structures observed in colloidal aggregates [25, 26, 27], and to predict others.
Our design strategy is validated by numerical calculations for several model systems and by ab-initio calculations for alkali metal clusters. The predictions of our theory are further confirmed by simulations of the growth of alkali and transition metal nanoparticles including up to four different elements. These simulations show that atoms naturally self-assemble into the predicted multi-shell icosahedral structures, including those with symmetry-breaking shell sequences. Although the applications presented below concern atomic clusters and nanoparticles, due to the general character of its basic assumptions, our theory can be applied to the design of aggregates of other particle types, such aggregates of colloidal particles and complex molecules of biological interest.
Results
@@
Here we develop our theoretical framework, whose starting point of is the well-known approach of Caspar and Klug (CK) [14, 16, 15], which was originally proposed for rationalizing and predicting the architecture of icosahedral viral capsids. Specifically, they developped a general method for the construction of individual icosahedral shells, which is based on cutting and folding leaflets from the two-dimensional hexagonal lattice and produces achiral and chiral arrangements of particles on the icosahedral surface.
In the CK theory, the triangulation number $T$ of an icosahedral shell is defined as the square length of the triangular edge, and is calculated as $T=h^{2}+k^{2}+hk$. Edge and radius of the shell are $\sqrt{T}$ and $\sin(2\leavevmode{}/5)\sqrt{T}$, respectively. Assigning one particle to each lattice point, the shell contains $210\,T+$ 2 particles (see Supplementary Note 1.1).
Shells are achiral or chiral depending on the angle $\theta_{\text{CK}}$ between the $h$-axis and the segment of the CK construction. Achiral shells correspond to segments with $\theta_{\text{CK}}=0^{\circ},30^{\circ}$ and 60. Segments on a coordinate axis correspond to the achiral shells described by Mackay (MC) in his work on the packing of equal spheres [13]. Shells built on segments on the diagonal ($h=k$, $\theta_{\text{CK}}=30^{\circ}$) are here called of Bergman (BG) type, since the smallest is the outer shell of the Bergman cluster [30, 38]. All other shells are chiral, with enantiomers symmetrically placed with respect to the diagonal. We remark that, in the CK theory, an icosahedral shell is uniquely determined by the segment endpoint $(h,k)$ in the hexagonal lattice; therefore, in the following, icosahedral shells will be denoted by their $(h,k)$.
@@
In general, paths can be drawn by allowing different increments at each step. Here we deal with the simplest generalization, which consists of choosing between $(h,k)\to(h,k+1)$ and $(h,k)\to(h+1,k)$ (other types of steps in the hexagonal plane will be discussed in the following). In this case, the number of shells in a path from the origin to a point $(h,k)$ is $i=h+k+1$. Such paths are inspired by the Mackay path of Figure 2a, in which one index is incremented at steps of one, whereas the other is kept constant and equal to zero. Allowing for different elementary steps in the hexagonal plane allows to build a wide variety of icosahedral structures, which retain the densely-packed character of the Mackay icosahedron.
@@
When assembling shells in physical systems, one must bear in mind that in the icosahedron the radius is shorter than the edge by $\sin(2\leavevmode{}/5)\approx 0.9511$, which has a direct effect on the packing of concentric shells. In the Mackay icosahedron of equal spheres [13], the distance between spheres in neighbouring shells is shorter by about $5\%$ than that between spheres in the same shell. Similar considerations hold for shells belonging to other chirality classes, assembled according to the path rules identified so far; in some cases, the difference between intra-shell and inter-shell nearest-neighbour distances is even larger. These considerations naturally lead to the idea of assembling shells in which particles in different shells have different sizes, which we better clarify below.
We consider the case of a core with $i$ shells, to which we add the outer shell $i+1$. Particles in the core and in the outer shell have different sizes. We define the size mismatch $\text{sm}_{i,i+1}=(d_{i+1}-d_{i})/d_{i}$, with $d_{i+1}$ and $d_{i}$ particle sizes in shells $i+1$ and $i$, respectively.
In the example of Figure 3a-b, the core is made of five MC shells, i.e. it is terminated by the $(0,4)$ shell. According to our path rules, the outer shell can be either the $(0,5)$ MC shell (same class) or the $(1,4)$ Ch1 shell (different class). We consider a simple model, in which all particles interact by the well-known Lennard-Jones (LJ) potential; specifically, the interaction energy is the same for all particles, the only difference being the equilibrium distance of pair interactions, which accounts for particles of different sizes (see the Methods section). In Figure 2a, we calculate the binding energy depending on the size mismatch, and determine the mismatch that minimizes the energy of the whole aggregate. For both MC and Ch1 shells this optimal mismatch is positive, i.e. it is favourable to have bigger particles in the outer shell than in the core, and it is larger for Ch1 than for MC shells.
The optimal mismatch can be estimated also by geometric packing arguments, which are discussed in details in Supplementary Note 1. Here we recall only the main assumption and give the final result. In order to evaluate the optimal mismatch between particles of shells $i$ and $i+1$ it is reasonable to impose that
@@
where $r_{i}$ and $r_{i+1}$ are the radii of the respective shells, $d_{i}$ and $d_{i+1}$ are the particle diameters in these shells. For atomic systems, these diameters are estimated from the nearest-neighbour distance of atoms in their crystal lattice. Recalling that the relation between the radii and the triangulation numbers of the shells
$r_{i}=d_{i}\sin\left(\frac{2\leavevmode{}}{5}\right)\sqrt{T_{i}}$ (2)
@@
$\text{sm}_{i,i+1}=\frac{2\sin\left(\frac{2\leavevmode{}}{5}\right)(1+\xi)\sqrt{T_{i}}+1}{2\sin\left(\frac{2\leavevmode{}}{5}\right)\sqrt{T_{i+1}}-1}-1.$ (3)
@@
In Figure 3c,d the results of Eq. (3) are compared to those for LJ and Morse clusters. Data in Figure 3c are obtained as in Figure 3a, but for different sizes of the Mackay core, whereas in Figure 3d we consider a cluster made of a Mackay core plus a Ch1 shell, to which we add a further shell of either Ch1 or Ch2 type. In all cases the agreement is good. Eq. (3) is thus a reliable guide for the semi-quantitative evaluation of the optimal mismatch, which demonstrates the key role of geometric factors in determining it.
@@
Our theory is now applied to the path-based design of icosahedral aggregates. In order to construct a multi-shell icosahedral aggregate, a path is drawn according to the rules in Figure 2 and, for each step, the optimal mismatch is estimated by Eq. (3). Then, particles of the appropriate sizes are associated with each shell.
An example is shown in Figure 4, where the path connecting BG shells through neighbouring chiral shells is considered (Figure 4a). This path alternates class-conserving and class-changing steps, spontaneously breaking mirror symmetries after the first BG shell. It produces the series of magic numbers
@@
with $b=4$ and $b=9$ for even and odd $i$, i.e. $N_{i}=1,13,45,117,239,431,...$ (see Supplementary Note 1.2).
The path of Fig. 4a is used to design multi-species alkali clusters. The values of size mismatch between alkali atoms, estimated from nearest-neighbour distances in bulk crystals, are in the range of the optimal values for this path (Figure 4b). In addition, these species present a weak tendency to mix and the bigger atoms have a smaller cohesive energy; this produces a general tendency for the bigger atoms to stay in the surface layers, which, as we have seen in the previous section, is exactly what is needed for stabilizing icosahedral aggregates.
The clusters in Figure 4c correspond to $i=3$ (Na13@K32, Na@${}_{13}@$Rb32) and to $i=4$ (Na@${}_{13}@$K@${}_{32}@$K72, Na@${}_{13}@$Rb@${}_{32}@$Rb72, Na@${}_{13}@$K@${}_{32}@$Rb72). In these clusters, the atomic species is always changed in class-changing steps, where a large mismatch is required, while it is changed or not in class-conserving steps, where the optimal mismatch is moderate, and therefore zero mismatch is expected to be acceptable.
The energetic stability of these clusters is verified in two ways: by full global optimization searches, using a semi-empirical force field, and by density functional theory (DFT) calculations on atomic pair exchanges (see Methods for details). Our global optimization searches find that the structures designed according to the path in Figure 4 are the lowest in energy for various binary and ternary alkali systems. Complete discussion and data are reported in Supplementary Note 4.1. The DFT data are reported in Figure 5 and Supplementary Tables 1-3 of Supplementary Note 2. In the case of CK shell sequences, exchanges of pairs of different atoms in adjacent shells produce energy increases. In contrast, if MC shells of these species are assembled in the same order, the resulting clusters are energetically unstable with respect to exchanges of atomic pairs, because the mismatch between species is too large for MC shells.
We have also checked the thermodynamic stability of MC@BG and MC@BG@Ch1 structures of Figure 4c, by performing molecular dynamics simulations in which the temperature is progressively increased up to melting. Our simulations reveal that the global minimum structures are highly stable, as they do not undergo any structural transformation until the cluster entirely melts (see Supplementary Note 4.2).
Our theory therefore provides the possibility of constructing icosahedra in systems where the traditional Mackay icosahedra would not be stable.
@@
We have demonstrated how to geometrically construct multi-shell icosahedra and verified their energetic stability in various systems. Another important point is to understand how they grow dynamically in physical processes. To this end, we have performed molecular dynamics (MD) growth simulations [44], in which atoms are deposited on pre-formed clusters. This type of simulations has been used to interpret previous nanoparticle growth experiments [24, 44]. The results are shown in Figure 6 and in Supplementary Figures 7-11 of Supplementary Note 3.1.
As a first case, we consider the growth of multi-element alkali clusters. In the previous section we have demonstrated the stability of non-trivial multi-shell structures, which comprise BG and chiral shells in the same aggregate. Here we check whether these structures can be grown in a physically realistic process. Here and in the following, growth temperatures are chosen to be significantly lower than the melting temperatures of the clusters.
@@
At this point, according to our rules, two non-equivalent steps are possible: the class-conserving step to $(1,3)$ and the class-changing step to $(2,2)$. These steps correspond to different optimal mismatches, i.e. 0.06 and 0.11, respectively, as estimated by Eq. (3). Since we are keeping on depositing Rb atoms (i.e. with zero mismatch), we expect growth to proceed by the $(1,3)$ step, which has the lower optimal mismatch. Also in this case, our prediction is verified by the growth simulation of Figure 6a, which further continues within the Ch1 class.
@@
Further confirmation of path rule predictivity is provided in Figure 6c-e. In Figure 6c, the size mismatch for the pairs AuNi, AuCo, AuFe, AgNi, AgCo, AgCu is compared with the optimal estimates obtained from Eq. (3) for the addition of one MC and one Ch1 shell on a Mackay core. The mismatch much better corresponds to Ch1 than to MC shells. For AuCo, AuFe, AuNi, AgNi and AgCo, Ch1 shells should grow on Mackay cores of 147 or 309 atoms ($k=3$ or $4$), while for AgCu the core should be larger, of 561 atoms ($k=5$). This is confirmed by MD simulations (Figure 6d-e) in which a Ch1 shell spontaneously forms when depositing Ag atoms on Ni147 or Cu561 Mackay cores. The growth continues within the Ch1 class if further Ag atoms are deposited. More results are presented and discussed in Supplementary Note 3.
We have verified that these structures are energetically stable. Icosahedra made of a Ni, Cu or Co Mackay core surrounded by one Ch1 shell of Ag atoms are the lowest in energy in several cases [39]. When adding a second Ch1 shell, the structures are not the lowest in energy any more, but they are in close competition with the global minima (see Supplementary Note 4.1). Anyway, when heated up by molecular dynamics simulations, these MC@Ch1@Ch1 clusters preserve their structure up to the melting temperature. Data are reported in Supplementary Note 4.2.
@@
We note that almost perfect shell-by-shell growth is achieved in all simulations due to the fast diffusion of deposited atoms on top of the close-packed shells [24].
Here we have shown and discussed the growth at some selected simulation temperatures. In Supplementary Note 3.4 we discuss the effect of temperature on the growth, showing that this type of icosahedral growth process is very robust, as it takes place in a wide range of temperatures.
@@
From there on, the growth proceeds in a different way from that of Figure 6e: a second shell of AM2 type is formed, followed by a shell of BG type, i.e. the class is changed at each step, even though atoms of the same type are deposited. This behaviour is due to the kinetic of the growth process. Specifically, it is due to fourfold adsorption sites on the surface of AM shells, which act as adatom traps and naturally lead to the formation of AM shells of higher class. The fourfold traps hinder the atomic mobility, causing a much rougher growth than in the case of Figure 6e. This point is discussed in detail in Supplementary Note 3.3. We note that, even though kinetic effects dominate the process, the growth proceeds step by step by incrementing only one shell index, as predicted by our rules.
@@
From the energetic point of view, also the generalized anti-Mackay structures may present notable stability. In fact, Ni147@Pd132 and Ni147@Pd132@Ag192 generalized anti-Mackay structures are the lowest in energy (see also Supplementary Note 4.1).
Discussion
Here we discuss the potential applications of the icosahedral structures described in this work, and we suggest possible extensions of our design strategy.
In many cases, the applications of nanoparticles stem from the structure of their surface. Our work shows that, to obtain a specific type of surface shell, it is necessary to build the correct sequence of shells in the inner part of the nanoparticles. This is possible if different elements (two or more) are used, with the appropriate mismatch. Applications of different kinds are also possible. For metal nanoparticles, it may be useful to protect a core (for example a magnetic core made of Ni, Co or Fe) by covering it with another metal. Our results indicate that for this purpose the size mismatch must be carefully chosen, depending also on the size of the magnetic core. If not, other asymmetric quasi-Janus structures with off-centered core would be much more stable (for example see ref. [46] and Supplementary Figure 15), and the core would be protected much less effectively.
In this work, we have treated specific examples related to metal clusters and nanoparticles. However, we believe that our construction can be relevant for other systems, e.g. colloidal aggregates. Colloids interact by potentials with very short range [47]. In Supplementary Note 1.6 we have shown that, for systems with short-range interaction, it is still possible to build stable structures according to our rules. However, the structures are less tolerant of deviations from the optimal size mismatch between shells, therefore being stable in a narrower range of mismatch values around the optimal one. This implies that shells must be carefully assembled to satisfy the constraints on mismatch values. In addition, we note that in some experiments on hard-sphere colloids, in which particle-particle attractive forces are negligible, icosahedral structures have been obtained in confined environments due to entropic effects [25, 26, 27]. These experiments were considering colloids of equal sizes and obtained aggregates with Mackay or anti-Mackay arrangements. Our results may be relevant also for these systems, since they may give a systematic guide on how to build those confined colloidal systems using colloids with different sizes. In fact, our approach, which is mostly based on general geometric considerations, is able to indicate the appropriate size mismatch between the spheres.
A natural generalization of the results presented here concerns in drawing paths that do not obey to the step rule $(h,k)\to(h,k+1)$ or $(h,k)\to(h+1,k)$. These paths are discussed in Supplementary Note 5, where we show that they are less likely to be physically relevant because they lead to poor matching between atoms in adjacent shells.